OpenStudy (anonymous):

simplify (sin^4 x - cos^4 x)/(sin x + cos x)

OpenStudy (tkhunny):

Have you considered factoring? Differences of Squares are wild!

OpenStudy (luigi0210):

^That. Do you know the formula?

OpenStudy (ranga):

\[ a^2 - b^2 = (a-b)(a+b)\]\[ a^4 - b^4 = (a^2)^2 - (b^2)^2 = (a^2-b^2)(a^2+b^2) = (a-b)(a+b)(a^2+b^2)\]

OpenStudy (anonymous):

thank you all

OpenStudy (anonymous):

i will try ot hammer it out and if i cant ill be back!

OpenStudy (anonymous):

\[\sin ^{3}x+sinxcos ^{2}x+-cosxsin ^{2}x+-\cos ^{3}x\]

OpenStudy (anonymous):

\[ \frac{\sin ^4(x)-\cos ^4(x)}{\sin (x)+\cos (x)}=\\\frac{(\sin (x)-\cos (x)) (\sin (x)+\cos (x)) \left(\sin ^2(x)+\cos ^2(x)\right)}{\sin (x)+\cos (x)}=\\\sin (x)-\cos (x) \]

OpenStudy (anonymous):

We used the fact that \[ \sin ^2(x)+\cos ^2(x)=1 \]

OpenStudy (anonymous):

ahhhh the cancellation and that pythagorean

OpenStudy (anonymous):

sinx-cosx, does that simplify?

OpenStudy (anonymous):

i guess thats it. i tried to find a theorem for that but couldnt